OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through the points (-3,-4) and (3,4).

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you need the slope first

OpenStudy (mathmath333):

\(y−y1=m(x−x1)\) \(m=\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{(x_1,y_1)=(-3,-4)~ and~ (x_2,y_2)=(3,4).}\end{align}\)

OpenStudy (anonymous):

\[m=\frac{ 4-(-4) }{ 3-(-3) }\]

OpenStudy (anonymous):

8/6

jimthompson5910 (jim_thompson5910):

always reduce fractions as much as possible

OpenStudy (anonymous):

4/3

jimthompson5910 (jim_thompson5910):

that is your slope

OpenStudy (anonymous):

ok so the equation is y=mx+b

OpenStudy (anonymous):

m=4/3

OpenStudy (anonymous):

but what is b

jimthompson5910 (jim_thompson5910):

they want the answer in point-slope form

jimthompson5910 (jim_thompson5910):

y=mx+b is slope intercept form

jimthompson5910 (jim_thompson5910):